# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ Appendix G Additional Experiments
Table 3: Main Comparison Results on Cora and Citeseer (with variance). We present the AUC scores for attack methods across different edge groups, now including mean $\pm$ variance. The best results are highlighted in bold, while the second-best results are underlined.
@@
Table 4: Main Comparison Results on Pubmed and LastFM-Asia (with variance). We present the AUC scores for attack methods across different edge groups, now including mean $\pm$ variance. The best results are highlighted in bold, while the second-best results are underlined.
Variance of Comparison Results.
@@
Due to space limitations, the comparison results in Table 1 in Section 6 do not include the standard variance from five repeated experiments. We now supplement all variance results in Table 3 and Table 4. We make the following observations regarding the stability of our proposed TrendAttack: Compared with the no-trend-feature counterparts MIA-GNN and StealLink, our TrendAttack-MIA and TrendAttack-SL exhibit smaller variances, indicating better stability. This highlights the effectiveness of the trend features, which not only improve attack performance on both edge groups but also reduce variance, enhancing model stability.
(a) Cora
(b) Citeseer
(c) Pubmed
Figure 7: Ablation study on the impact of trend feature order. Overall attack AUC as a function of the trend feature order (0–3) for three unlearn methods across three datasets.
Impact of Trend Feature Orders.
We investigate the influence of trend feature orders on the performance of the proposed TrendAttack framework. Following the experimental setup described in Section 6.2, we employ the most effective variant, TrendAttack-SL, for this ablation study. Specifically, we vary the trend feature order $k$ (see Algorithm 2) to assess the trade-off between incorporating additional neighborhood information and achieving high attack performance. The results are presented in Figure 7, from which we make the following observations:
(i) Incorporating trend features of orders 1, 2, or 3 significantly improves attack performance compared to the 0-th order (which corresponds to a degenerate form of TrendAttack that reduces to a simple StealLink attack). This highlights the effectiveness of our proposed trend feature design.
(ii) The attack performance remains relatively stable across orders 1 to 3, indicating that the method is robust to the choice of trend order. Notably, lower-order features (e.g., order 1) still have strong performance while requiring less auxiliary neighborhood information, making them more practical in real-world inversion attack scenarios.
Appendix H Impact Statement
@@
In this work, we propose a novel membership inference attack targeting unlearned GNNs, revealing critical privacy vulnerabilities in existing graph unlearning methods. Our goal is to raise awareness about privacy protection in Web services and to inspire future research on privacy-preserving graph machine learning. While our method introduces a new attack strategy, all experiments are conducted on publicly available benchmark datasets. Given the gap between this pioneering study and real-world deployment scenarios, we do not foresee significant negative societal implications from this work.

### v2

@@ Appendix G Additional Experiments
Table 3. Main Comparison Results on Cora and Citeseer (with variance). We present the AUC scores for attack methods across different edge groups, now including mean $\pm$ variance. The best results are highlighted in bold, while the second-best results are underlined.
@@
Table 4. Main Comparison Results on Pubmed and LastFM-Asia (with variance). We present the AUC scores for attack methods across different edge groups, now including mean $\pm$ variance. The best results are highlighted in bold, while the second-best results are underlined.
@@
Variance of Comparison Results. Due to space limitations, the comparison results in Table 1 in Section 6 do not include the standard variance from five repeated experiments. We now supplement all variance results in Table 3 and Table 4. We make the following observations regarding the stability of our proposed TrendAttack: Compared with the no-trend-feature counterparts MIA-GNN and StealLink, our TrendAttack-MIA and TrendAttack-SL exhibit smaller variances, indicating better stability. This highlights the effectiveness of the trend features, which not only improve attack performance on both edge groups but also reduce variance, enhancing model stability.
(a) Cora results.
(b) Pubmed results.
(c) LastFM-Asia results.
Figure 8. Ablation study on the impact of unlearning ratio.
(a) GIF results.
(b) CEU results.
(c) GA results.
Figure 9. Transferability of TrendAttack across different datasets on Cora, Citeseer, and Pubmed. Each heatmap shows attack AUC when the shadow dataset (rows) and attack dataset (columns) differ.
Dataset Transferability. We evaluate the ability of TrendAttack-SL to transfer across datasets using a GCN backbone. For each unlearning method (GIF, CEU, GA), we train the attack on one dataset (shadow dataset) and apply it to a different dataset (attack dataset). Figures 9 (a)-(c) present the results on Cora, Citeseer, and Pubmed. Overall, the attack maintains strong performance even when the shadow and attack datasets differ, with only slight decreases in some cases. This demonstrates the stability and generalization ability of TrendAttack-SL across datasets, indicating that the attack does not strongly depend on having the same dataset for training the shadow model.
Table 5. Comparison of our attack setting with previous unlearning inversion attacks. Our setting targets GNNs and assumes no access to the original model $f(\cdot;\bm{\theta}_{\mathrm{orig}})$, which is unlikely to be shared by responsible model owners due to its memorization of sensitive contents. Although we use a shadow dataset, our assumption of only black-box access to the unlearned model $f(\cdot;\bm{\theta}_{\mathrm{un}})$ makes our setting one of the strictest and most practical in the unlearning inversion literature.
Method Access to $f(\cdot;\bm{\theta}_{\mathrm{orig}})$ Access to $f(\cdot;\bm{\theta}_{\mathrm{un}})$ Shadow Dataset Victim Model Attacker’s Goal
ReconAttack (bertran2024reconstruction) White-box White-box No Linear regression Unlearned features
UnlearnInv-Feature (hu2024learn) White-box White-box No DNNs Unlearned features
UnlearnInv-Label (hu2024learn) Black-box Black-box No DNNs Unlearned labels
Ours No Black-box Yes GNNs Unlearned edges
Appendix H Discussions
In this section, we discuss the rationale of our attack settings and present potential future directions for potential defence strategies of our work.
H.1. Comparison to Previous Unlearning Inversion Attacks
Compared to prior unlearning inversion attacks (bertran2024reconstruction; hu2024learn) (see Table 5), our attack is conducted under one of the strictest threat models in the literature. The main difference lies in model access assumptions: both ReconAttack (bertran2024reconstruction) and UnlearnInv (hu2024learn) adopt a two-model setting, where the attacker can access both the original model $f_{\mathrm{orig}}$ (before unlearning) and the unlearned model $f_{\mathrm{un}}$ (after unlearning). In contrast, our method uses a one-model setting, where the attacker can only access the unlearned model $f_{\mathrm{un}}$. This difference is important, as in realistic deployments the pre-unlearning model, which memorizes sensitive data, is rarely released. Therefore, our attack setting is fundamentally different and reflects practical scenarios. Detailed comparisons are as follows.
Difference to ReconAttack (bertran2024reconstruction). This work targets linear regression models with different loss functions, while our attack addresses a much more complex setting involving GNNs. Our method explicitly considers how unlearning effects propagate over graph structures and tackles challenges from deep, non-linear architectures, which go beyond simple regression models. Moreover, (bertran2024reconstruction) assumes white-box access to both the original and unlearned models’ parameters (see “…the parameter difference between the two models…” on page 2 of (bertran2024reconstruction)), whereas our method only requires black-box access to the unlearned model’s outputs.
Difference to UnlearnInv (hu2024learn). This work presents two settings. The feature attack (Section 3.2.1 of (hu2024learn)) assumes white-box access to both models and aims to recover training features. The label attack (Section 3.2.2 of (hu2024learn)) assumes black-box access to both models and infers training labels. Our TrendAttack, by contrast, requires only black-box access to $f_{\mathrm{un}}$ and uses a shadow dataset. Instead of recovering features or labels, we focus on reconstructing unlearned edges, an attack goal uniquely important for graph-structured data.
H.2. The Necessity of Shadow Datasets
In this work, we first train the attack model on the shadow dataset $\mathcal{G}^{\mathrm{sha}}$, and then transfer the trained attack model to the target dataset and its corresponding victim model (i.e., the unlearned GNN trained on the target dataset). An alternative strategy is to discard the shadow dataset and instead train a link prediction model directly on the target dataset $\mathcal{G}_{\mathrm{un}}$, which is a more intuitive way to recover unlearned edges from the cleaned graph. However, due to API query limitations imposed by many model-serving platforms and online social networks (e.g., see the query limitation policies of Twitter [^2] https://developer.x.com/en/docs/x-api and TikTok [^3] https://developers.tiktok.com/doc/research-api-faq), attackers can typically collect only a small number of user profiles and social connections within a given time window. As a result, it is difficult to gather sufficient training data directly on the target dataset, making such direct link prediction approaches less practical in real-world settings. Thus, our setting of training the attack model on a shadow dataset with a similar distribution to the target dataset (e.g., Facebook in the US $\rightarrow$ Facebook in the UK) provides access to more training data and is therefore more practical under such widespread query limitations.
H.3. Potential Defence Strategies
Leveraging Existing Defences for MIAs. Since our proposed attack is a significantly more challenging variant of traditional membership inference attacks (MIAs), defences developed for MIAs may still offer protection. For example, training differentially private GNNs, as suggested in (hu2024learn), could help mitigate our attack.
Strengthening Unlearning Mechanisms. Our attack exploits the residual effects left by imperfect unlearning. Increasing the unlearning magnitude (i.e., allocating a stronger privacy budget) could help reduce such effects. However, finding a good balance between model utility and privacy remains a challenge.
Post-Unlearning Mitigation Strategies. It may also be helpful to adopt mitigation strategies for general-purpose unlearning inversion attacks on Euclidean (i.e., non-graph) data. Techniques such as parameter pruning or fine-tuning have been shown to reduce privacy leakage (hu2024learn), and could be applied to reduce the effectiveness of inversion attacks like ours.
Ethical Consideration
@@ H.3. Potential Defence Strategies
In this work, we propose a novel privacy attack targeting unlearned GNNs, revealing critical privacy vulnerabilities in existing graph unlearning methods. Our goal is to raise awareness about privacy protection in Web services and to inspire future research on privacy-preserving graph machine learning. While our method introduces a new attack strategy, all experiments are conducted on publicly available benchmark datasets. Given the gap between this pioneering study and real-world deployment scenarios, we do not foresee significant negative societal implications from this work.